OpenStudy (anonymous):

Suppose there is a toy which takes either 'a' or 'b' as input. Total number of inputs=n. If two consecutive inputs are 'a' the toy crashes. What is the total number of ways in which the toy does NOT crash?

OpenStudy (anonymous):

Looks like an offsetted fibbonacci sequence. f(1) = 2 ; a b f(2) = 3 ; ab ba bb f(3) = 5 ; abb aba bab bbb bba f(4) = 8 ; abbb abba abab babb baba bbba bbbb bbab f(5) = 13 etc. In general: f(1) = 2 , f(2) = 3, f(n) = f(n-1) + f(n-2) ; n>=3 See http://www.mathsisfun.com/numbers/fibonacci-sequence.html

OpenStudy (anonymous):

So basically you're saying that f(n) for some n is the some of the first n fibonacci numbers?

OpenStudy (anonymous):

rather it is the nth term of the fibonacci sequence?

OpenStudy (anonymous):

Ok thanks a lot sir/ma'm....

OpenStudy (anonymous):

m=1 ; total ways not to crash toy is fib(4) = 2 m=2 ; total ways not to crash toy is fib(5) = 3 : : m=n ; total ways not to crash toy is fib(n+3) So, for n inputs, the total number of ways not to crash the toy is the (n+3)'th term of the fibonacci sequence

OpenStudy (anonymous):

Actually - in the fibonacci series, terms are counted from 0: fib(0)=0 , fib(1)=1, fib(2)=1, fib(3)=2, fib(4)=3, fib(5)=5, fib(6)=8, fib(7)=13, ... So, for n inputs, the total number of ways not to crash the toy is the (n+2)'th term of the fibonacci sequence (0,1,1,2,3,5,8,13,...) counting the terms from 0. Then you can also use this formula to calculate the answer for any n: \[fib(n)=\frac{\phi^{n}-(1-\phi)^{n}}{\sqrt{5}}\] where phi is the golden ratio , approximately: \[\phi=\frac{1+\sqrt{5}}{2}\] See: http://www.mathsisfun.com/numbers/fibonacci-sequence.html http://www.mathsisfun.com/numbers/golden-ratio.html